# MM-LLMs: Recent Advances in MultiModal Large Language Models

### 25 Jan 2024 v2

@@ 6 Future Directions
We can enhance the MM-LLMs’ strength from the following four key avenues: (1) Expanding Modalities: Current MM-LLMs typically support the following modalities: image, video, audio, 3D, and text. However, the real world involves a broader range of modalities. Extending MM-LLMs to accommodate additional modalities (e.g., web pages, heat maps, and figures&tables) will increase the model’s versatility, making it more universally applicable; (2) Diversifying LLMs: Incorporating various types and sizes of LLMs provides practitioners with the flexibility to select the most appropriate one based on their specific requirements; (3) Improving MM IT Dataset Quality: Current MM IT dataset have ample room for improvement and expansion. Diversifying the range of instructions can enhance the effectiveness of MM-LLMs in understanding and executing user commands. (4) Strengthening MM Generation Capabilities: Most current MM-LLMs are predominantly oriented towards MM understanding. Although some models have incorporated MM generation capabilities, the quality of generated responses may be constrained by the capacities of the LDMs. Exploring the integration of retrieval-based approaches Asai et al. (2023) holds significant promise in complementing the generative process, potentially enhancing the overall performance of the model.
@@
Existing benchmarks might not adequately challenge the capabilities of MM-LLMs, given that many datasets have previously appeared to varying degrees in the PT or IT sets. This implies that the models may have learned these tasks during training. Moreover, current benchmarks predominantly concentrate on the VL sub-field. Thus, it is crucial for the development of MM-LLMs to construct a more challenging, larger-scale benchmark that includes more modalities and uses a unified evaluation standard. Concurrently, benchmarks can be tailored to assess the MM-LLMs’ proficiency in practical applications. For instance, the introduction of GOAT-Bench Lin et al. (2024) aims to evaluate various MM-LLMs’ capacity to discern and respond to nuanced aspects of social abuse presented in memes.
@@
To deploy MM-LLMs on resource-constrained platforms and achieve optimal performance meanwhile, such as low-power mobile and IoT devices, lightweight implementations are of paramount importance. A notable advancement in this realm is MobileVLM Chu et al. (2023a). This approach strategically downscales LLaMA, allowing for seamless off-the-shelf deployment. MobileVLM further introduces a Lightweight Downsample Projector, consisting of fewer than million parameters, contributing to improved computational speed. Nevertheless, this avenue necessitates additional exploration for further advancements in development.
@@
The embodied intelligence aims to replicate human-like perception and interaction with the surroundings by effectively understanding the environment, recognizing pertinent objects, assessing their spatial relationships, and devising a comprehensive task plan Firoozi et al. (2023). Embodied AI tasks, such as embodied planning, embodied visual question answering, and embodied control, equips robots to autonomously implement extended plans by leveraging real-time observations. Some typical work in this area is PaLM-E Driess et al. (2023) and EmbodiedGPT Mu et al. (2023). PaLM-E introduces a multi-embodiment agent through the training of a MM-LLM. Beyond functioning solely as an embodied decision maker, PaLM-E also demonstrates proficiency in handling general VL tasks. EmbodiedGPT introduces an economically efficient method characterized through a CoT approach, enhancing the capability of embodied agents to engage with the real world and establishing a closed loop that connects high-level planning with low-level control. While MM-LLM-based Embodied Intelligence has made advancements in integrating with robots, further exploration is needed to enhance the autonomy of robots.
Continual IT
In practical applications, MM-LLMs are expected to adapt to new MM tasks for supporting additional functionalities. Nevertheless, current MM-LLMs remain static and are unable to adjust to continuously emerging requirements. Therefore, an approach is needed to make the model flexible enough to efficiently and continually leverage emerging data, while avoiding the substantial cost of retraining MM-LLMs. This aligns with the principles of continual learning, where models are designed to incrementaly learn new tasks similar to human learning. Continual IT aims to continuously fine-tune MM-LLMs for new MM tasks while maintaining superior performance on tasks learned during the original MM IT stage. It introduces two primary challenges: (1) catastrophic forgetting, where models forget previous knowledge when learning new tasks Robins (1995); McCloskey and Cohen (1989); Goodfellow et al. (2013); Zhang et al. (2023d, c, b); Zheng et al. (2023a), and (2) negative forward transfer, indicating that the performance of unseen tasks is declined when learning new ones Zheng et al. (2024); Dong et al. (2023b, a). Recently, He et al. established a benchmark to facilitate the development of continual IT for MM-LLMs. Despite these advancements, there is still a significant opportunity and room for improvement in developing better methods to address the challenges of catastrophic forgetting and negative forward transfer.
@@ Appendix A Related Surveys
We have comprehensively covered nearly all MM-LLMs over the past year, including not only understanding models but also generative models. Our coverage extends beyond VL modalities to encompass various modes such as audio and 3D;
@@
Appendix B Mainstream PEFT Methods
@@
Appendix C Commonly Used LLMs
The commonly used LLM Backbones in existing MM-LLMs research are as follows:
@@ Appendix C Commonly Used LLMs
ChatGLM222https://github.com/THUDM/ChatGLM-6B is a Chinese-English bilingual dialogue model, optimized by an auto-regressive mask infilling objective. It is based on the GLM Du et al. (2022b); Zeng et al. (2022a) architecture, optimized for Chinese question answering and dialogues.
@@
UL2 Tay et al. (2022) is an encoder-decoder model trained utilizing a mixture of denoisers objectives, surpassing T5 on numerous benchmarks.
@@
Qwen Bai et al. (2023a) is trained on large-scale and diverse datasets, with a primary focus on Chinese and English. It employs SFT and RLHF techniques for alignment, resulting in dialogue models like Qwen-Chat.
@@
Chinchilla Hoffmann et al. (2022) is a causal decoder, trained on extensive text data. It posits that model size should double for every doubling of training tokens.
@@
PaLM Chowdhery et al. (2023) is a causal decoder structure with parallel attention and feed-forward layers, enabling training speeds up to times faster. Notable changes contain RoPE embeddings, SwiGLU activation, multi-query attention, and etc.
@@
LLaMA Touvron et al. (2023a) comprises decoder-only models with efficient causal attention.
@@
LLaMA-2 Touvron et al. (2023b) focuses on fine-tuning a superior and safer LLaMA-2-Chat model for conversation generation, incorporating 40% more training data with grouped-query attention and a larger context length.
@@
Vicuna Chiang et al. (2023) is a model built on top of LLaMA, utilizing user dialogue data obtained from ShareGPT.com and trained by SFT.
Appendix D SOTA MM-LLMs (continued)
@@
(20) LLaVA-1.5 Liu et al. (2023d) reports simple modifications to the LLaVA framework, including applying an MLP projection and introducing VQA data tailored for academic tasks, along with simple response formatting prompts. These adjustments result in enhanced capabilities for MM understanding.
@@
(21) MiniGPT-v2 Chen et al. (2023c) is an MM-LLM designed as a unified interface for diverse VL multi-task learning. To create a single model proficient in handling multiple VL tasks, identifiers are incorporated for each task during both training and inference. This facilitates clear task distinction, ultimately enhancing learning efficiency.
@@
(22) CogVLM Wang et al. (2023) is an open-source MM-LLM that bridges the gap between modalities via a trainable visual expert module within the attention and feedforward layers. This allows for a deep fusion of MM features without compromising performance on NLP downstream tasks.
@@
(23) DRESS Chen et al. (2023h) introduces a method using natural language feedback to enhance alignment with human preferences. DRESS extends the conditional reinforcement learning algorithm to integrate non-differentiable natural language feedback, training the model to generate appropriate responses based on feedback.
@@
(24) X-InstructBLIP Panagopoulou et al. (2023) introduces a cross-modal framework with instruction-aware representations, scalable enough to empower LLMs to handle diverse tasks across multiple modalities, including image/video, audio, and 3D. Notably, it achieves this without the need for modality-specific PT.
@@
(25) CoDi-2 Tang et al. (2023a) is a MM generation model excelling in modality-interleaved instruction following, in-context generation, and user-model interaction by multi-turn conversations. It enhances CoDi Tang et al. (2023b) to process intricate modality-interleaved inputs and instructions, generating latent features autoregressively.
@@
(26) VILA Lin et al. (2023) outperforms in vision tasks and shows remarkable reasoning ability while maintaining text-only capabilities. It achieves this by harnessing the full capabilities of LLM learning, using the interleaved attributes of image-text pairs, and implementing meticulous text data re-blending.
@@
| Episodic WebLI Chen et al. (2023g) | Image | 400M | 400M | 400M |
@@
| MMC4 (Interleaved) Zhu et al. (2023b) | Image | 571M | 43B | 101.2M (Instances) |
@@
| AISHELL-2 Chen et al. (2023b) | Audio | – | – | 128K |
@@
| VSDial-CN Chen et al. (2023b) | Image, Audio | 120K (Image), 1.2M(Audio) | 120K | 1.2M |
@@
| StableLLaVA Li et al. (2023f) | SFT | I+TT | SD Rombach et al. (2022) | Auto.+Manu. | ✗ | 126K/–/– | 1 | 126K |
| LLaVA’s IT Zhang et al. (2023f) | SFT | I+TT | MS-COCO | Auto. | ✔ | 81K/–/– | 2.29 | 150K |
@@
| LLaVAR Zhang et al. (2023f) | SFT | I+TT | MS-COCO, CC3M, LAION | LLaVA+Auto. | ✔ | 20K/–/– | 2.27 | 174K |
| ShareGPT4V Chen et al. (2023e) | SFT | I+TT | LCS, COCO, SAM, TextCaps, WikiArt | Auto.+Manu. | ✗ | 100K/–/– | – | – |
| DRESS’s IT Chen et al. (2023h) | SFT | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 193K/–/– | 4 | – |
| VideoChat’s IT Li et al. (2023d) | SFT | V+TT | WebVid | Auto. | ✔ | –/8K/– | 1.82 | 11K |
@@
| MIT Li et al. (2023e) | SFT | I/V/B+TT | Multiple | Auto.+Manu. | ✗ | –/–/– | 1 | 2.4M |
@@
| mPLUG-DocOwl’s IT Ye et al. (2023) | SFT | I/Tab/Web+TT | Multiple | Inherit | ✗ | – | – | – |
@@
| DRESS’s IT Chen et al. (2023h) | RLHF | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 33K/–/– | 4 | – |
Appendix E VL Benchmarks
The VL benchmarks presented in Table 2 include OKVQA Schwenk et al. (2022), IconVQA Lu et al. (2021), VQA Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQA: ScienceQA-IMG Lu et al. (2022), VQA: TextVQA Singh et al. (2019), POPE Li et al. (2023g), MME: MME Perception Fu et al. (2023), MME: MME Cognition Fu et al. (2023), MMB: MMBenchmark Liu et al. (2023f), MMB: MMBench-Chinese Liu et al. (2023f), SEED: SEED-Bench (Image) Li et al. (2023b), LLaVA: LLaVA-Bench (In-the-Wild) Liu et al. (2023a), MM-Vet Yu et al. (2023), QBench Wu et al. (2023b), HM: HatefulMemes Kiela et al. (2020), and VSR Liu et al. (2023a).

### 20 Feb 2024 v4

@@ 6 Future Directions
We can enhance the MM-LLMs’ strength from the following four key avenues: (1) Expanding Modalities: Current MM-LLMs mainly support the following modalities: image, video, audio, 3D, and text. However, the real world involves a broader range of modalities. Extending MM-LLMs to accommodate additional modalities (e.g., web pages, heat maps, and figures&tables) will increase the model’s versatility, making it more universally applicable; (2) Diversifying LLMs: Incorporating various types and sizes of LLMs provides practitioners with the flexibility to select the most appropriate one based on their specific requirements; (3) Improving MM IT Dataset Quality: Current MM IT datasets have ample room for improvement and expansion. Diversifying the range of instructions can enhance the effectiveness of MM-LLMs in understanding and executing user commands; (4) Strengthening MM Generation Capabilities: Most current MM-LLMs are predominantly oriented towards MM understanding. Although some models have incorporated MM generation capabilities, the quality of generated responses may be constrained by the capacities of the LDMs. Exploring the integration of retrieval-based approaches Asai et al. (2023); Gao et al. (2023a) holds significant promise in complementing the generative process, potentially enhancing the overall performance of the model.
@@
Existing benchmarks might not adequately challenge the capabilities of MM-LLMs, given that many datasets have previously appeared to varying degrees in the PT or IT sets. This implies that the models may have learned these tasks during training. Moreover, current benchmarks predominantly concentrate on the VL sub-field. Thus, it is crucial for the development of MM-LLMs to construct a more challenging, larger-scale benchmark that includes more modalities and uses a unified evaluation standard. For instance, GOAT-Bench Lin et al. (2024b) is introduced to assess the capability of various MM-LLMs in discerning and responding to nuanced aspects of social abuse depicted in memes. MathVista Lu et al. (2024) evaluates the math reasoning ability of MM-LLMs within visual contexts. Moreover, MMMU Yue et al. (2023) and CMMMU Zhang et al. (2024) have respectively introduced English and Chinese versions of the massive multi-discipline MM understanding and reasoning benchmark for expert artificial general intelligence. Fan et al. have also challenged MM-LLMs with multipanel VQA. BenchLMM Cai et al. (2023) benchmarks the cross-style visual capability of MM-LLMs. Additionally, Liu et al. have conducted an in-depth study on the optical character recognition capabilities of MM-LLMs.
@@
To deploy MM-LLMs on resource-constrained platforms and achieve optimal performance meanwhile, such as low-power mobile and IoT devices, lightweight implementations are of paramount importance. A notable advancement in this realm is MobileVLM Chu et al. (2023a). This approach strategically downscales LLaMA, allowing for seamless off-the-shelf deployment. MobileVLM further introduces a lightweight downsample projector, consisting of fewer than million parameters, contributing to improved computational speed. Recently, there have been many similar studies on lightweighting MM-LLMs, achieving efficient computation and inference with comparable performance or minimal loss, including TinyGPT-V Yuan et al. (2023b), Vary-toy Wei et al. (2024), Mobile-Agent Wang et al. (2024b), MoE-LLaVA Lin et al. (2024a), and MobileVLM V2 Chu et al. (2024). Nevertheless, this avenue necessitates additional exploration for further advancements in development.
@@
The embodied intelligence aims to replicate human-like perception and interaction with the surroundings by effectively understanding the environment, recognizing pertinent objects, assessing their spatial relationships, and devising a comprehensive task plan Firoozi et al. (2023). Embodied AI tasks, such as embodied planning, embodied visual question answering, and embodied control, equip robots to autonomously implement extended plans by leveraging real-time observations. Some typical works in this area are PaLM-E Driess et al. (2023) and EmbodiedGPT Mu et al. (2023). PaLM-E introduces a multi-embodiment agent through the training of a MM-LLM. Beyond functioning solely as an embodied decision maker, PaLM-E also demonstrates proficiency in handling general VL tasks. EmbodiedGPT introduces an economically efficient method characterized by a CoT approach, enhancing the capability of embodied agents to engage with the real world and establishing a closed loop that connects high-level planning with low-level control. While MM-LLM-based Embodied Intelligence has made advancements in integrating with robots, further exploration is needed to enhance the autonomy of robots.
Continual Learning
Due to the large training costs associated with their massive scale, MM-LLMs are not amenable to frequent re-training. However, updates are necessary to endow MM-LLMs with new skills and keep them up-to-date with rapidly evolving human knowledge Wu et al. (2024). Thus, Continual Learning (CL) is needed to make the model flexible enough to efficiently and continually leverage emerging data while avoiding the substantial cost of retraining MM-LLMs. CL for MM-LLMs can be classified into two stages: continual PT and continual IT. Recently, a continual MM IT benchmark has been proposed to continuously fine-tune MM-LLMs for new MM tasks while maintaining superior performance on tasks learned during the original MM IT stage He et al. (2023). It introduces two primary challenges: (1) catastrophic forgetting, where models forget previous knowledge when learning new tasks Robins (1995); McCloskey and Cohen (1989); Goodfellow et al. (2013); Zhang et al. (2023d, c, b); Zheng et al. (2023a), and (2) negative forward transfer, indicating that the performance of unseen tasks declines when learning new ones Zheng et al. (2024)
Mitigating Hallucination
Hallucinations entail generating textual descriptions of nonexistent objects without visual cues, which manifest in diverse categories Liu et al. (2024a) such as misjudgments and inaccuracies in descriptions. The origins of these hallucinations are multifaceted Liu et al. (2024a), including biases and annotation errors in training data. Additionally, Skip Han et al. (2024) highlights semantic drift biases associated with paragraph separators, which can induce hallucinations when deliberately inserted. Current methods to mitigate these hallucinations involve leveraging self-feedback as visual cues Lee et al. (2023). However, challenges persist, necessitating nuanced discernment between accurate and hallucinatory outputs, as well as advancements in training methodologies to enhance output reliability.
@@ Appendix A Related Surveys
We have comprehensively covered nearly all MM-LLMs over the past year, totaling around or more, including not only understanding models but also generative models. Our coverage extends beyond VL modalities to encompass various modes such as audio and 3D point cloud;
@@
Appendix B Modality Encoder
In the following, we provide a brief introduction to some mainstream modality encoders.
B.1 Visual Modality
NFNet-F6
ViT
CLIP ViT
Radford et al. (2021) builds connections between text and images, comprising a ViT and a text encoder. With a vast amount of text-image pairs, it optimizes ViT by contrastive learning, treating paired text and images as positive samples and others as negative ones.
Eva-CLIP ViT
Fang et al. (2023) stabilizes the training and optimization process of the massive CLIP, offering new directions in expanding and accelerating the expensive training of MM base models.
B.2 Audio Modality
C-Former
HuBERT
BEATs
Chen et al. (2023g) is an iterative audio pre-training framework designed to learn Bidirectional Encoder representations from Audio Transformers.
Appendix C Mainstream PEFT Methods
@@
LayerNorm tuning
Zhao et al. (2024) presents an efficient strategy to transform LLMs into MM-LLMs, which tunes LayerNorm in attention block yielding strong MM performance compared with full parameter finetuning or LoRA.
In addition to the aforementioned PEFT methods, there are several others, including P-tuning Liu et al. (2022), P-tuning v2 Liu et al. (2021a), AdaptBias Fu et al. (2022), Compacter Karimi Mahabadi et al. (2021), AdapterFormer Chen et al. (2022a), XTuner Contributors (2023), P-LoRA Dong et al. (2024b), MoLE Chen et al. (2024), and Delta-LoRA Zi et al. (2023).
Appendix D Representative LLMs
The representative LLM Backbones in existing MM-LLMs research are as follows:
Flan-T5
ChatGLM
@@ Appendix D Representative LLMs
is a Chinese-English bilingual dialogue model,222https://github.com/THUDM/ChatGLM-6B optimized by an auto-regressive mask infilling objective. It is based on the GLM Du et al. (2022b); Zeng et al. (2022a) architecture, optimized for Chinese question answering and dialogues.
InternLM
Team (2023) is a multilingual trillion-parameter foundation model trained on over a trillion tokens of data. Based on this foundation, the model utilizes high-quality human-annotated dialogue data combined with RLHF to respond to complex instructions during human interactions, exhibiting responses that align with human ethics and values.
UL2
@@
Tay et al. (2022) is an encoder-decoder model trained utilizing a mixture of denoisers objectives, surpassing T5 on numerous benchmarks.
Qwen
@@
Bai et al. (2023a) is trained on large-scale and diverse datasets, with a primary focus on Chinese and English. It employs SFT and RLHF techniques for alignment, resulting in dialogue models like Qwen-Chat.
Chinchilla
@@
Hoffmann et al. (2022) is a causal decoder, trained on extensive text data. It posits that model size should double for every doubling of training tokens.
OPT
PaLM
@@
Chowdhery et al. (2023) is a causal decoder structure with parallel attention and feed-forward layers, enabling training speeds up to times faster. Notable changes contain RoPE embeddings, SwiGLU activation, multi-query attention, and etc.
LLaMA
@@
Touvron et al. (2023a) comprises decoder-only models with efficient causal attention.
LLaMA-2
@@
Touvron et al. (2023b) focuses on fine-tuning a superior and safer LLaMA-2-Chat model for conversation generation, incorporating 40% more training data with grouped-query attention and a larger context length.
Vicuna
@@
Chiang et al. (2023) is a model built on top of LLaMA, utilizing user dialogue data obtained from ShareGPT.com and trained by SFT.
Appendix E SOTA MM-LLMs
In the following, we will provide a brief introduction to the core contributions of some representative MM-LLMs.
Flamingo
Alayrac et al. (2022) represents a series of Visual Language (VL) Models designed for processing interleaved visual data and text, generating free-form text as the output.
BLIP-2
Li et al. (2023e) introduces a more resource-efficient framework, comprising the lightweight Q-Former to bridge modality gaps and the utilization of frozen LLMs. Leveraging LLMs, BLIP-2 can be guided for zero-shot image-to-text generation using natural language prompts.
LLaVA
Liu et al. (2023e) pioneers the transfer of IT techniques to the MM domain. Addressing data scarcity, LLaVA introduces a novel open-source MM instruction-following dataset created using ChatGPT/GPT-4, alongside the MM instruction-following benchmark, LLaVA-Bench.
MiniGPT-4
Zhu et al. (2023a) proposes a streamlined approach where training only one linear layer aligns the pre-trained vision encoder with the LLM. This efficient method enables the replication of the exhibited capabilities of GPT-4.
mPLUG-Owl
Ye et al. (2023b) presents a novel modularized training framework for MM-LLMs, incorporating the visual context. To assess different models’ performance in MM tasks, the framework includes an instructional evaluation dataset called OwlEval.
X-LLM
Chen et al. (2023b) is expanded to various modalities, including audio, and demonstrates strong scalability. Leveraging the language transferability of the Q-Former, X-LLM is successfully applied in the context of Sino-Tibetan Chinese.
VideoChat
Li et al. (2023f) pioneers an efficient chat-centric MM-LLM for video understanding dialogue, setting standards for future research in this domain and offering protocols for both academia and industry.
InstructBLIP
Dai et al. (2023) is trained based on the pre-trained BLIP-2 model, updating only the Q-Former during MM IT. By introducing instruction-aware visual feature extraction and corresponding instructions, the model enables the extraction of flexible and diverse features.
PandaGPT
Su et al. (2023) is a pioneering general-purpose model with the capability to comprehend and act upon instructions across different modalities: text, image/video, audio, thermal, depth, and inertial measurement units.
(PaLI-X
Chen et al. (2023h) is trained using mixed VL objectives and unimodal objectives, including prefix completion and masked-token completion. This approach proves effective for both downstream task results and achieving the Pareto frontier in the fine-tuning setting.
Video-LLaMA
Zhang et al. (2023e) introduces a multi-branch cross-modal PT framework, enabling LLMs to simultaneously process the vision and audio content of a given video while engaging in conversations with humans. This framework aligns vision with language as well as audio with language.
Video-ChatGPT
Maaz et al. (2023) is a model specifically designed for video conversations, capable of generating discussions about videos by integrating spatiotemporal vision representations.
Shikra
Chen et al. (2023e) introduces a simple and unified pre-trained MM-LLM tailored for Referential Dialogue, a task involving discussions about regions and objects in images. This model demonstrates commendable generalization ability, effectively handling unseen settings.
DLP
Jian et al. (2023) proposes the P-Former to predict the ideal prompt, trained on a dataset of single-modal sentences. This showcases the feasibility of single-modal training to enhance MM learning.
BuboGPT
Zhao et al. (2023d) is a model constructed by learning a shared semantic space for a comprehensive understanding of MM content. It explores fine-grained relationships among different modalities such as image, text, and audio.
ChatSpot
Zhao et al. (2023b) introduces a simple yet potent method for finely adjusting precise referring instructions for MM-LLM, facilitating fine-grained interactions. The incorporation of precise referring instructions, consisting of image- and region-level instructions, enhances the integration of multi-grained VL task descriptions.
Qwen-VL
Bai et al. (2023b) is a multi-lingual MM-LLM that supports both English and Chinese. Qwen-VL also allows the input of multiple images during the training phase, improving its ability to understand the vision context.
NExT-GPT
Wu et al. (2023d) is an end-to-end, general-purpose any-to-any MM-LLM that supports the free input and output of image, video, audio, and text. It employs a lightweight alignment strategy, utilizing LLM-centric alignment in the encoding phase and instruction-following alignment in the decoding phase.
MiniGPT-5
Zheng et al. (2023b) is an MM-LLM integrated with inversion to generative vokens and integration with Stable Diffusion. It excels in performing interleaved VL outputs for MM generation. The inclusion of classifier-free guidance during the training phase enhances the quality of generation.
LLaVA-1.5
@@ Appendix E SOTA MM-LLMs
Liu et al. (2023d) reports simple modifications to the LLaVA framework, including applying an MLP projection and introducing VQA data tailored for academic tasks, along with simple response formatting prompts. These adjustments result in enhanced capabilities for MM understanding.
MiniGPT-v2
@@
Chen et al. (2023d) is an MM-LLM designed as a unified interface for diverse VL multi-task learning. To create a single model proficient in handling multiple VL tasks, identifiers are incorporated for each task during both training and inference. This facilitates clear task distinction, ultimately enhancing learning efficiency.
CogVLM
@@
Wang et al. (2023b) is an open-source MM-LLM that bridges the gap between modalities via a trainable visual expert module within the attention and feedforward layers. This allows for a deep fusion of MM features without compromising performance on NLP downstream tasks.
DRESS
@@
Chen et al. (2023i) introduces a method using natural language feedback to enhance alignment with human preferences. DRESS extends the conditional reinforcement learning algorithm to integrate non-differentiable natural language feedback, training the model to generate appropriate responses based on feedback.
X-InstructBLIP
@@
Panagopoulou et al. (2023) introduces a cross-modal framework with instruction-aware representations, scalable enough to empower LLMs to handle diverse tasks across multiple modalities, including image/video, audio, and 3D. Notably, it achieves this without the need for modality-specific PT.
CoDi-2
@@
Tang et al. (2023b) is a MM generation model excelling in modality-interleaved instruction following, in-context generation, and user-model interaction by multi-turn conversations. It enhances CoDi Tang et al. (2023c) to process intricate modality-interleaved inputs and instructions, generating latent features autoregressively.
VILA
@@
Lin et al. (2023) outperforms in vision tasks and shows remarkable reasoning ability while maintaining text-only capabilities. It achieves this by harnessing the full capabilities of LLM learning, using the interleaved attributes of image-text pairs, and implementing meticulous text data re-blending.
@@
| Episodic WebLI Chen et al. (2023h) | Image | 400M | 400M | 400M |
@@
| MMC4 (Interleaved) Zhu et al. (2023c) | Image | 571M | 43B | 101.2M (Instances) |
| Obelics (Interleaved) Laurençon et al. (2023) | Image | 353M | 115M | 141M (Instances) |
@@
| AISHELL-1 Chen et al. (2023b) | Audio | – | – | 128K |
@@
| VSDial-CN Ni et al. (2024) | Image, Audio | 120K (Image), 1.2M(Audio) | 120K | 1.2M |
@@
| StableLLaVA Li et al. (2023i) | SFT | I+TT | SD Rombach et al. (2022) | Auto.+Manu. | ✗ | 126K/–/– | 1 | 126K |
| LLaVA’s IT Zhang et al. (2023h) | SFT | I+TT | MS-COCO | Auto. | ✔ | 81K/–/– | 2.29 | 150K |
@@
| LLaVAR’s IT Zhang et al. (2023h) | SFT | I+TT | MS-COCO, CC3M, LAION | LLaVA+Auto. | ✔ | 20K/–/– | 2.27 | 174K |
| ShareGPT4V’s IT Chen et al. (2023f) | SFT | I+TT | LCS, COCO, SAM, TextCaps, WikiArt | Auto.+Manu. | ✗ | 100K/–/– | – | – |
| DRESS’s IT Chen et al. (2023i) | SFT | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 193K/–/– | 4 | – |
| VideoChat’s IT Li et al. (2023f) | SFT | V+TT | WebVid | Auto. | ✔ | –/8K/– | 1.82 | 11K |
@@
| MIT Li et al. (2023h) | SFT | I/V/B+TT | Multiple | Auto.+Manu. | ✗ | –/–/– | 1 | 2.4M |
@@
| mPLUG-DocOwl’s IT Ye et al. (2023b) | SFT | I/Tab/Web+TT | Multiple | Inherit | ✗ | – | – | – |
@@
| Osprey’s IT Yuan et al. (2023a) | SFT | I+TT | MS-COCO, RefCOCO, RefCOCO+, LLaVA’s IT etc. (fine-grained region-text dataset) | Auto.+Manu. | ✔ | –/–/– | 4 | 724K |
| LLaVA-RLHF Sun et al. (2023b) | RLHF | I+TT | Collected human preference | Manu. | ✗ | –/–/– | – | 10K |
| DRESS’s IT Chen et al. (2023i) | RLHF | I+TT | LLaVA’s IT, VLSafe | Auto.+Manu. | ✔ | 33K/–/– | 4 | – |
| RLHF-V’s IT Yu et al. (2023b) | RLHF | I+TT | Collected human preference | Manu. | ✗ | –/–/– | – | 1.4K |
| VLFeedback Li et al. (2023g) | RLHF | I+TT | Responses generated by MM-LLMs | Auto. | ✗ | –/–/– | – | 80K |
| RTVLM Li et al. (2024a) | RLHF | I+TT | New question-image pairs based on publicly available images or originally diffusion-generated images Gallegos et al. (2023) | Auto.+Manu. | ✗ | –/–/– | – | 5K |
| VLGuard’s IT Zong et al. (2024) | RLHF | I+TT | Source image data from various datasets | Auto. | ✗ | 3K/–/– | – | 3K |
| MMViG Yan et al. (2024) | RLHF | I+TT | MS-COCO | Manu. | ✗ | 16K/–/– | – | 16K |
Appendix F VL Benchmarks
The VL benchmarks presented in Table 2 include OKVQA Schwenk et al. (2022), IconVQA Lu et al. (2021), VQA Goyal et al. (2017), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQA: ScienceQA-IMG Lu et al. (2022), VQA: TextVQA Singh et al. (2019), POPE Li et al. (2023k), MME: MME Perception Fu et al. (2023), MME: MME Cognition Fu et al. (2023), MMB: MMBenchmark Liu et al. (2023g), MMB: MMBench-Chinese Liu et al. (2023g), SEED: SEED-Bench (Image) Li et al. (2023c), LLaVA: LLaVA-Bench (In-the-Wild) Liu et al. (2023a), MM-Vet Yu et al. (2023c), QBench Wu et al. (2023b), HM: HatefulMemes Kiela et al. (2020), and VSR Liu et al. (2023a).